# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ 1. $(i)$
    If the barcode ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{1}}})=\{[b^{I}_{l_{1}},d^{I}_{l_{1}}]\}$ is not empty and satisfies $\left\lVert d^{I}_{l_{1}}-b^{I}_{l_{1}}\right\rVert_{\infty}>\left\lVert\overrightarrow{u}\right\rVert_{\infty}$, then the barcode ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{2}}})$ is not empty as well, and
@@ 1. $(ii)$
    If the barcodes ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{1}}})$ and ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{2}}})$ are not empty, then one has
Proof.
Item $(i)$. Since $|(d^{I}_{l_{1}})_{i}-(b^{I}_{l_{1}})_{i}|=\left\lVert d^{I}_{l_{1}}-b^{I}_{l_{1}}\right\rVert_{\infty}>\left\lVert\overrightarrow{u}\right\rVert_{\infty}$ for any index $i\in\llbracket 1,n\rrbracket$, it follows that $b^{I}_{l_{1}}\leq_{n}b^{I}_{l_{1}}+\overrightarrow{u}\leq_{n}d^{I}_{l_{1}}$. Thus $b^{I}_{l_{1}}+\overrightarrow{u}$ must belong to ${\rm supp}(I)$ since $I$ is an interval. Hence, since $b^{I}_{l_{1}}+\overrightarrow{u}\in l_{2}$, one has ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{2}}})={\rm supp}(I)\cap l_{2}\neq\varnothing$.
Item $(ii)$. If one of the endpoints is infinite, the result holds trivially as the other endpoint has to be infinite too, so we now assume that the endpoints of the bars are all finite. Without loss of generality, assume that $l_{2}=l_{1}+\overrightarrow{u}$ where $\overrightarrow{u}$ is positive. Now, since both $d_{l_{2}}^{I}$ and $d_{l_{1}}^{I}+\overrightarrow{u}$ belong to $l_{2}$, they are comparable, so one has either $d_{l_{2}}^{I}>_{n}d_{l_{1}}^{I}+\overrightarrow{u}$ or $d_{l_{2}}^{I}\leq_{n}d_{l_{1}}^{I}+\overrightarrow{u}$. However, the first possibility would lead to $d_{l_{2}}^{I}>_{n}d_{l_{1}}^{I}+\overrightarrow{u}>_{n}d_{l_{1}}^{I}$, hence $d_{l_{1}}^{I}$ and $d_{l_{2}}^{I}$ would be (strictly) comparable in $\mathbb{R}^{n}$, which contradicts Lemma 3.9. Thus, one must have $d_{l_{2}}^{I}\leq_{n}d_{l_{1}}^{I}+\overrightarrow{u}$. Furthermore, and using the exact same arguments, $d^{I}_{l_{2}}-\overrightarrow{u}+\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\boldsymbol{1}$ is on $l_{1}$, and one must have $d_{l_{2}}^{I}-\overrightarrow{u}+\left\lVert\overrightarrow{u}\right\rVert_{\infty}\cdot\boldsymbol{1}\geq_{n}d_{l_{1}}^{I}$. Finally, by combining the two previous inequalities, one has
Endpoint location.
The next definition and result show that endpoints of an interval module must be located in the vicinity of the other endpoints of the module that are close to it—more precisely, in their rectangle hull. This will be useful in the proof of Proposition 3.4; in particular, we will use this result to characterize the positions of endpoints of any given diagonal line $l$ solely from the endpoints of the lines of the grid $L$ that are close to $l$.
Definition 3.11.
Let $S\subseteq\mathbb{R}^{n}$. The rectangle hull of $S$, denoted by $\mathrm{recthull}[S]$, is defined as the smallest rectangle containing $S$
where $(\wedge S)_{i}:=\min_{s\in S}s_{i}$ and $(\vee S)_{i}:=\max_{s\in S}s_{i}$.
Lemma 3.12 (Endpoints bound).
@@
Proof.
Figure 5: Illustration of $H_{j},d,l,d^{j}$.
Since $d^{j}$ and $d$ belong to $H_{j}$, they have the same $j$-th coordinate: $d^{j}_{j}=d_{j}$. Moreover, both $d^{j}$ and $\pi_{l}(d)$ belong to the diagonal line $l$, hence they are comparable, and $\|d^{j}-\pi_{l}(d)\|_{\infty}=|(d^{j}-\pi_{l}(d))_{i}|$ for any $i\in\llbracket 1,n\rrbracket$. Then, one has $\|d^{j}-d\|_{\infty}\leq\|d^{j}-\pi_{l}(d)\|_{\infty}+\|\pi_{l}(d)-d\|_{\infty}=|(d^{j}-\pi_{l}(d))_{j}|+\|\pi_{l}(d)-d\|_{\infty}=|(d-\pi_{l}(d))_{j}|+\|\pi_{l}(d)-d\|_{\infty}\leq 2\|\pi_{l}(d)-d\|_{\infty}\leq\delta$. Let $d^{+}=d^{j}+\delta\sum_{j\in\mathcal{J}^{\prime\prime}}e_{j}\text{ and }d^{-}=d^{j}-\delta\sum_{j\in\mathcal{J}^{\prime}}e_{j}$, where
$\mathcal{J}^{\prime}=\left\{i\in\llbracket 1,n\rrbracket\backslash\{j\}:d_{i}<d^{j}_{i}\right\}\text{ and }\mathcal{J}^{\prime\prime}=\left\{i\in\llbracket 1,n\rrbracket\backslash\{j\}:d_{i}>d^{j}_{i}\right\}.$
By construction, one has $d^{-}\leq d\leq d^{+}\in H_{j}$ and
$\|d^{+}-d\|_{\infty},\|d^{-}-d\|_{\infty}\leq\delta.$ (4)
Proof of inequalities. We now show that $\overline{d}_{j}\geq_{n}d_{j}\geq_{n}\underline{d}_{j}$. We start with the second inequality. Since $d^{+}$ and $\underline{d}$ are one the same diagonal line, they are comparable. Furthermore, if one had $d^{+}<_{n}\underline{d}$ by contradiction, then the induced rectangle $R_{d,\underline{d}}$ would not be flat since $d\leq_{n}d^{+}<_{n}\underline{d}$, which would contradict Lemma 3.9. As a consequence, $d^{+}\geq_{n}\underline{d}$. Taking the $j$-th coordinate yields $d_{j}=d^{+}_{j}\geq\underline{d}_{j}$. The first inequality holds using the same arguments.
This proof applies straightforwardly to birthpoints by symmetry. ∎
Using Lemma 3.10, one can generalize Lemma 3.12 above to the case where some lines in $L_{x,\delta}$ have an empty intersection with ${\rm supp}(I)$, and then define a common location for all endpoints that belong to the convex hull of the same $L$-surrounding set, as we do in the following proposition.
Proposition 3.13.
Let $I$ be an interval module. Let $K$ be a hyperrectangle and $L=L_{\delta}(K^{\delta})$ be a $\delta$-grid of lines. Let $l\in L$ such that $|L_{l}|=2^{n-1}$ and assume that ${\rm conv}(L_{l})\cap L[I]$ (resp. ${\rm conv}(L_{l})\cap U[I]$) is not empty. Then, there exists a set $B_{l}$ (resp. $D_{l}$) such that for any $x\in{\rm conv}(L_{l})\cap L[I]$ (resp. ${\rm conv}(L_{l})\cap U[I]$), one has either $\|b_{x}^{I}-d_{x}^{I}\|_{\infty}\leq\delta$, or $x\in B_{l}$ (resp. $D_{l}$), where $B_{l}$ (resp. $D_{l}$) is a rectangular set in $\mathbb{R}^{n}$ that can be constructed from the birthpoints $(b_{l^{\prime}}^{I})_{l^{\prime}\in L_{l}}$ (resp. deathpoints $(d_{l^{\prime}}^{I})_{l^{\prime}\in L_{l}}$). Moreover, one has that
and similarly for $D_{l}$.
Proof.
We first construct $B_{l}$ and $D_{l}$, and then we will show items (1) and (2).
Definition of $B_{l},D_{l}$. Let first assume that $x$ is in the interior of ${\rm conv}(L_{l})$, that we denote with ${\kern 0.0pt{\rm conv}(L_{l})}^{\mathrm{o}}$. Note that if there is a line $l_{0}$ that is $\delta$-comparable to $l_{x}$, and such that ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l_{0}}})=\varnothing$, then by Lemma 3.10 $(i)$, one immediately has $\|b_{x}^{I}-d_{x}^{I}\|_{\infty}\leq\delta$. Hence, we now assume that the barcodes along any line that is $\delta$-comparable to $l_{x}$ is not empty, which means that the hypotheses of Lemma 3.12 are satisfied for $x$. Now, remark that since $L$ is a grid, if one is able to find a line $l^{\prime}$ in $L$ whose intersections with hyperplanes associated to the canonical axes of $\mathbb{R}^{n}$ are $\delta$-close to $x$, then, since $x$ is in the interior of an $L$-surrounding set $L_{l}$, $l^{\prime}$ must belong to that surrounding set $L_{l}$ as well. More formally, one has that, for any line $l^{\prime}\in L$,
This ensures (from Equation 4) that the lines of $L$ associated to $\tilde{B}^{I}_{x,\delta}$ and $\tilde{D}^{I}_{x,\delta}$ are all included in $L_{l}$ for any $x\in{\kern 0.0pt{\rm conv}(L_{l})}^{\mathrm{o}}$, and thus that we can safely define
Note that $B_{l}$ and $D_{l}$ depend only on the endpoints of the lines in $L_{l}$ and that $d_{x}^{I}\in D_{l}$ and $b_{x}^{I}\in B_{l}$ for any $x\in{\kern 0.0pt{\rm conv}(L_{l})}^{\mathrm{o}}$ by Lemma 3.12. Furthermore, if $x$ is in the closure of ${\rm conv}(L_{l})$, the previous statements still hold since $D_{l}$ and $B_{l}$ are closed sets. We now show that $B_{l}$ and $D_{l}$ satisfy Equation (5).
Proof of Equation (5). By applying Lemma 3.12 and its proof for arbitrary dimension $j$ to all $x\in{\rm conv}(L_{l})\cap U[I]$, there exist deathpoints $\overline{d}^{j}$ and $\underline{d}_{j}$ that satisfy $(\overline{d}^{j})_{j}=\sup_{d\in D_{l}}d_{j}$ and $(\underline{d}_{j})_{j}=\inf_{d\in D_{l}}d_{j}$ and $(\underline{d}_{j})_{j}\leq x_{j}\leq(\overline{d}^{j})_{j}$ for all $x\in{\rm conv}(L_{l})\cap U[I]$. Moreover, these points are located on lines in $L_{l}$ that are $\delta$-consecutive by definition. Thus, applying Lemma 3.10 $(ii)$ repeatedly on these pairs of lines for all dimensions, we end up with $D_{l}$ having a diagonal smaller than $\delta$. The same goes for birthpoints.
3.2.3 Proof of Proposition 3.4
Proof.
Let ${\mathbb{M}}=\bigoplus_{i\in\mathcal{I}}I_{i}$ and $\tilde{\mathbb{M}}_{\delta}=\bigoplus_{i\in\tilde{\mathcal{I}}}\tilde{I}_{i}$ be the interval decompositions of ${\mathbb{M}}$ and $\tilde{\mathbb{M}}_{\delta}$, with induced exact matchings $\sigma_{{\mathbb{M}}}$ and $\sigma_{\tilde{\mathbb{M}}_{\delta}}$ respectively. In order to upper bound the bottleneck distance ${d_{\rm B}}({\mathbb{M}},\tilde{\mathbb{M}}_{\delta})$, one can upper bound the interleaving distance ${d_{\rm I}}(I_{i},\tilde{I}_{\nu(i)})$ for any index $i\in\mathcal{I}$. Let $I$ and $\tilde{I}$ be two such intervals (we drop the index $i$ to alleviate notations). Since $I$ and $\tilde{I}$ are interval modules, and thus indicator modules, the morphisms $\varphi^{(\delta)}_{I\to\tilde{I}}\colon I\rightarrow\tilde{I}(\boldsymbol{\delta})$ and $\varphi^{(\delta)}_{\tilde{I}\to I}\colon\tilde{I}\rightarrow I(\boldsymbol{\delta})$ are well-defined. We thus need to show that these morphisms commute, i.e., that they induce a $\delta$-interleaving. Hence, we first show that
$\left(\varphi^{(\delta)}_{\tilde{I}\to I}\right)_{x+\boldsymbol{\delta}}\circ\left(\varphi^{(\delta)}_{I\to\tilde{I}}\right)_{x}=\left(\varphi^{(2\delta)}_{I}\right)_{x},$ (6)
for any $x\in K$.
If $x\in l$ for some line $l\in L$, Equation (6) is satisfied from ${\rm supp}(I)\cap l={\rm supp}(\tilde{I})\cap l$, which itself comes from the fact that $\tilde{\mathbb{M}}_{\delta}$ has the same $L$-fibered barcode than ${\mathbb{M}}$. Hence, we assume in the following that $x\not\in\cup_{l\in L}l$. Furthermore, if $x\not\in{\rm supp}(I)$ or $x+2\boldsymbol{\delta}\not\in{\rm supp}(I)$, then Equation (6) is trivially satisfied. Hence, we also assume $x,x+2\boldsymbol{\delta}\in{\rm supp}(I)$. This means that $b^{I}_{x}$ and $d^{I}_{x}$ are well-defined, and that $\left(\varphi^{(2\delta)}_{I}\right)_{x}\cong\mathrm{id}_{k\to k}$. Thus we only have to show that $\tilde{I}_{x+\boldsymbol{\delta}}\cong k$, i.e., $x+\boldsymbol{\delta}\in{\rm supp}(\tilde{I})$.
As $L=L_{\delta}(K^{\delta})$ is the $\delta$-grid of $K^{\delta}$ and $x\in K$, let $l\in L$ be a line such that $x\in{\rm conv}(L_{l})$ and let $l_{x}\subseteq{\rm conv}(L_{l})$ be the diagonal line passing through $x$. Now, as $R_{x,x+\boldsymbol{\delta}}\subseteq{\rm supp}(I)$, Lemma 3.10 $(i)$ ensures that ${\mathcal{B}}({\left.\kern-1.2ptI\vphantom{\big{|}}\right|_{l}})\neq\varnothing$ for any line $l\in L$ that is $\delta$-comparable to $l_{x}$; and the same holds for $\tilde{I}$ since ${\mathcal{F}}{\mathcal{B}}(I)_{L}={\mathcal{F}}{\mathcal{B}}(\tilde{I})_{L}$. Using Proposition 3.13 on both $I$ and $\tilde{I}$, there exist two sets $B_{l}$ and $D_{l}$ such that $d_{x}^{I},d_{x}^{\tilde{I}}\in D_{l}$ and $b_{x}^{I},b_{x}^{\tilde{I}}\in B_{l}$, with the segments $l_{x}\cap B_{l}$ and $l_{x}\cap D_{l}$ having length at most $\delta$. Since one also has
$b_{x}^{I}\leq x\leq x+2\boldsymbol{\delta}\leq d_{x}^{I},$
and $\left\lVert d_{x}^{I}-d_{x}^{\tilde{I}}\right\rVert_{\infty},\left\lVert b_{x}^{I}-b_{x}^{\tilde{I}}\right\rVert_{\infty}\leq\delta$, one finally has $b_{x}^{\tilde{I}}\leq x+\boldsymbol{\delta}\leq d_{x}^{\tilde{I}},$ which concludes that $x+\boldsymbol{\delta}\in{\rm supp}(\tilde{I})$. ∎
4 Computing stable candidate decompositions with MMA
In this section, we first introduce the general problem of computing candidate decompositions (see Definition 3.3) in Section 4.1, and we introduce our algorithm MMA for computing instances of candidate decompositions on any multi-parameter persistence module in arbitrary dimension, i.e., with arbitrary number of filtrations, in Section 4.2. In particular, one does not have to assume that the underlying module is decomposable in order to run MMA. When the underlying module is interval decomposable, we show that our algorithm MMA can actually recover it exactly in Section 4.3. Finally, even when the underlying module is not necessarily interval decomposable, we also show that the candidate decompositions computed by MMA are still stable in Section 4.4.
4.1 Motivation
The goal of this section is to frame the general question of practically computing candidate decompositions of a multi-parameter persistence module from its fibered barcode. There are many ways of doing so, but the most natural ones are not necessarily the easiest computable ones. For the sake of simplicity, let us leave the problem of finding proper matching functions aside for now (which we will discuss in more details in Section 5), and assume that the underlying module is a single interval module ${\mathbb{M}}=I$. Since interval modules are characterized by their supports, the goal is to recover ${\rm supp}(I)$. Moreover, if $I$ is discretely presented, only the facets and critical points of ${\rm supp}(I)$ need to be captured or approximated. There are many different ways, for a given interval module $I$, to define candidate critical points, that we call corners, using the endpoints of its fibered barcode, e.g., by using the minimum and maximum of consecutive endpoint coordinates. Hence, it is natural to define candidate decompositions (or candidate approximations in this case, since there is just one summand) $\tilde{I}$ with model selection, i.e., by minimizing some penalty cost $\mathrm{pen}\colon S\to\mathbb{R}_{+}$, where $S$ is the set of discretely presented interval modules having the same fibered barcode as $I$, or a subset thereof. See Figure 6 for examples of sets $S$ and corresponding candidate approximations. This penalty would forbid, e.g., overly complicated approximations that have a lot of corners. For instance, minimizing the penalty
$\mathrm{pen}:\tilde{I}\mapsto\#\text{corners of }{\rm supp}(\tilde{I}).$ (7)
would provide a sparse approximation of $I$. Actually, when one assumes that the underlying interval module $I$ is discretely presented with facets that are large enough with respect to the family of lines $L$ of the fibered barcode, the target $I$ minimizes penalty (7). Indeed, as all the facets of $I$ are detected by some endpoints of the fibered barcode by assumption, any candidate approximation $\tilde{I}$ of $I$ has at least the same number of facets than $I$, i.e., $\mathrm{pen}(I)\leq\mathrm{pen}(\tilde{I})$ for any candidate approximation $\tilde{I}$.
@@ 4.1 Motivation
Figure 6: Example of candidates for a $2$-interval module $I$ with support in $\mathbb{R}^{2}$. (Left) Given the $L$-fibered barcode of $I$, where $L$ is the family of the four black lines, we want to approximate $I$ with an element of $S$, i.e., an interval module with the same fibered barcode. (Middle) When one further constrains the set $S$ by asking to have at most one corner between two consecutive endpoints of the fibered barcode, the whole set $S$ can be computed explicitly. (Right) The set $S$ can also be described as the set of intervals which have to go through the blue path, and which can arbitrarily choose between the red or green path at three different locations. Hence, the cardinality of $S$ is $2^{3}$.
@@
For interval modules, $S$ is generally a set of cardinal $c^{d}$, where $c$ is the number of candidate corners between birthpoints or deathpoints, and $d$ is the number of corners. For instance, in Figure 6, one has $n=2$, $c=2$ and $d=3$. Unfortunately, $c$ is of the order of $2^{n-1}$, and thus grows exponentially with the dimension $n$, and $d$ is difficult to control in practice, since it heavily depends on the number of lines in the fibered barcode and the regularity of the underlying interval module $I$. Minimizing a penalty over $S$ is thus practical only for low dimension $n$ and small number of lines in the fibered barcode. Hence, our algorithm MMA presented in Section 4.2 does not use penalty minimization, but is rather defined with natural and simple corner choices.
Remark 4.1.
@@
Note also that there are cases when the corner choices are canonical. For instance, any $2$-multipersistence module $M$ with transition maps $\varphi_{\cdot}^{\cdot}$ that are weakly exact, i.e., that satisfy, for any $x\leq y$
@@
is rectangle decomposable [BLO22]. Hence, a canonical approximation of a summand $I$ of ${\mathbb{M}}$ is given by the interval module whose support is the rectangle with corners $(\min_{l}(b^{I}_{l})_{1},\min_{l}(b^{I}_{l})_{2})$ and $(\max_{l}(b^{I}_{l})_{1},\max_{l}(b^{I}_{l})_{2})$, where $l$ goes through the family of lines $L$ of the fibered barcode.
4.2 MMA algorithm for computing candidate decompositions
In this section, we introduce MMA: a fast algorithm for computing $\delta$-candidate decompositions. The pseudo-code for MMA is provided in Algorithm 1. Roughly speaking, given a multi-parameter persistence module ${\mathbb{M}}$, an approximation parameter $\delta>0$, a $\delta$-grid of lines $L=L_{\delta}(K^{\delta})$ where $K$ contains the multigraded Betti numbers of ${\mathbb{M}}$, and a matching function $\sigma$ that commutes with the exact matching function $\sigma_{\mathbb{M}}$ induced by ${\mathbb{M}}$ (see Section 5 for a discussion about how to find such matching functions), Algorithm 1 works in three steps:

### v4

@@ 1. $(i)$
    $\tilde{\mathbb{M}}_{\delta}$ is interval decomposable: $\tilde{\mathbb{M}}_{\delta}=\bigoplus_{i\in{\tilde{\mathcal{I}}}}{k^{\tilde{I}_{i}}}$, and
@@ 2. $(ii)$
    ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\subseteq{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}_{\delta}\vphantom{\big|}\right|_{l}}\right)$ for any $l\in L$, i.e., the $L$-fibered barcode of $\mathbb{M}$, seen as a multiset of segments in $\mathbb{R}^{n}$, is included in the one of $\tilde{\mathbb{M}}_{\delta}$.
Clearly, a candidate decomposition can be a rough descriptor of $\mathbb{M}$, as the bars in its fibered barcode can be arbitrarily large. Hence, we identify a more stable subfamily of candidate decompositions:
  1. $(iii)$
    If ${d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}_{\delta}\vphantom{\big|}\right|_{l}}\right)\right)\leq 2\delta$ for any diagonal line $l$ (not only those that belong to $L$), then $\tilde{\mathbb{M}}_{\delta}$ is called an approximate decomposition of $\mathbb{M}$.
The reason we focus on preserving the diagonal fibered barcodes (instead of controlling, e.g., the rank invariant or the interleaving distance to $\mathbb{M}$) is because of the impossibility for general multi-parameter persistence modules to build a decomposition into indicator modules that is consistent with the rank invariant (see (46, Section 10.2.3)). Note however that this is still stronger than preserving the Hilbert function, i.e., the pointwise dimension of the module.
Remark 5 (Non-diagonal lines).
Extending the definition of our candidate decompositions to grids of non-diagonal lines (i.e., with direction vector different than $\boldsymbol{e}_{\Delta}$) is straightforward, and is completely equivalent to rescaling the filtrations. Using such non-diagonal grids will however produce less stable descriptors, as the interleaving distance (Definition 4) is based on the diagonal direction.
Remark 6.
One can check that the $\delta$-grid of lines of associated to $K^{\delta}$ used in Definition 19 is $\delta$-regularly distributed and $\delta$-fills $K$.
Finally, we introduce the definition of matching functions. Such functions play a key role in our MMA algorithm for computing candidate decompositions.
Definition 20 (Matching function).
Let $\mathbb{M}$ be a f.p. $n$-parameter persistence module, and $l,l^{\prime}\subseteq\mathbb{R}^{n}$ be two positive lines. A map $\sigma$ between the persistence barcodes:
$\sigma\colon{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\rightarrow{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right)\cup\left\{\varnothing\right\}$
is called an ($\mathbb{M}$-)matching function between $l$ and $l^{\prime}$ if the restriction of $\sigma$ to $\sigma^{-1}({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right))$ is injective. In other words, $\sigma$ is a partial bijection (Definition 5) between the two barcodes, seen as multisets of intervals.
Definition 21 (Induced matching functions).
If $\mathbb{M}=\bigoplus_{i\in\mathcal{I}}k^{I_{i}}$ is a f.p. interval decomposable module, then, for any positive line $l$, the bars of any barcode ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\cong\bigoplus_{i\in\mathcal{I}}{\left.\kern-1.2ptk^{{I_{i}}}\vphantom{\big|}\right|_{l}}$ can be indexed using $\mathcal{I}$ (by also counting empty bars). Thus, given any two positive lines $l_{1},l_{2}$, one can match the bars $\left.\kern-1.2ptk^{I_{i}}\vphantom{\big|}\right|_{l_{1}}$} and $\left.\kern-1.2ptk^{{I_{i}}}\vphantom{\big|}\right|_{l_{2}}$} together so that matched bars correspond to the same underlying interval summand of $\mathbb{M}$. In that case, the corresponding matching function $\sigma_{\mathbb{M}}$ is referred to as induced from $\mathbb{M}$.
3.2 Motivation for the MMA algorithm
Our MMA algorithm can be roughly described as a method that constructs interval summands based on families of bars (coming from the fibered barcode) that have been matched together using some matching function. The goal of this section is to frame the general question of practically computing candidate decompositions of a multi-parameter persistence module from its fibered barcode and a matching function. There are many ways of doing so, but the most natural ones are not necessarily the easiest computable ones. For the sake of simplicity, let us leave the problem of finding proper matching functions aside for now (which we will discuss in more details in Section 6), and assume that the underlying module is a single interval module $\mathbb{M}=\mathbb{I}$. Since interval modules are characterized by their supports, the goal is to recover ${\rm supp}(\mathbb{I})$. Moreover, if $\mathbb{I}$ is discretely presented, only the facets and critical points (i.e., points where several facets intersect) of ${\rm supp}(\mathbb{I})$ need to be captured or approximated. There are many different ways, for a given interval module $\mathbb{I}$, to define candidate critical points, that we call corners, using the endpoints of its fibered barcode, e.g., by using the minimum and maximum of consecutive endpoint coordinates. Hence, it is natural to find a candidate decomposition (or candidate interval in this case, since there is just one interval summand) $\tilde{\mathbb{I}}$ with model selection, i.e., by minimizing some penalty cost $\mathrm{pen}\colon S\to\mathbb{R}_{+}$, where $S$ is the set of discretely presented interval modules having the same fibered barcode than $\mathbb{I}$, or a subset thereof. See Figure 5 for examples of sets $S$ and corresponding candidate intervals. This penalty would forbid, e.g., overly complicated intervals that have lots of corners. For instance, minimizing the penalty:
@@ 3.2 Motivation for the MMA algorithm
[…]
@@
Figure 5: Example of candidate decompositions for a 22-interval module $\mathbb{I}$ with support in $\mathbb{R}^{2}$. (Left) Given the $L$-fibered barcode of $\mathbb{I}$, where $L$ is the family of the four black lines, we want to approximate $\mathbb{I}$ with an element of $S$, i.e., an interval module with the same fibered barcode. (Middle) When one further constrains the set $S$ by asking to have at most one corner between two consecutive endpoints of the fibered barcode, the whole set $S$ can be computed explicitly. (Right) The set $S$ can also be described as the set of intervals which have to go through the blue path, and which can arbitrarily choose between the red or green path at three different locations. Hence, the cardinality of $S$ is $2^{3}$.
@@
For interval modules, $S$ is generally a set of cardinal $c^{d}$, where $c$ is the number of candidate corners between birthpoints or deathpoints, and $d$ is the number of corners. For instance, in Figure 5, one has $n=2$, $c=2$ and $d=3$. Unfortunately, $c$ is of the order of $2^{n-1}$, and thus grows exponentially with the dimension $n$, and $d$ is difficult to control in practice, since it heavily depends on the number of lines in the fibered barcode and the regularity of the underlying interval module $\mathbb{I}$. Minimizing a penalty over $S$ is thus practical only for low dimension $n$ and small number of lines in the fibered barcode. Hence, our algorithm MMA presented in Section 3.3 does not use penalty minimization, but is rather defined with natural and simple corner choices.
Remark 7.
@@
Note also that there are cases when the corner choices are canonical. For instance, any $2$-persistence module $\mathbb{M}$ with transition maps $\varphi_{\cdot}^{\cdot}$ that are weakly exact, i.e., that satisfy, for any $x\leq y$:
@@
is rectangle decomposable (10). Hence, a canonical approximation of a summand $\mathbb{I}$ of $\mathbb{M}$ is given by the interval module whose support is the rectangle with corners $(\min_{l}(b^{I}_{l})_{1},\min_{l}(b^{I}_{l})_{2})$ and $(\max_{l}(b^{I}_{l})_{1},\max_{l}(b^{I}_{l})_{2})$, where $l$ goes through the family of lines $L$ of the fibered barcode.
3.3 The MMA algorithm for computing candidate decompositions
In this section, we introduce MMA: a fast algorithm for computing $\delta$-candidate decompositions. The pseudo-code for MMA is provided in Algorithm 1. Roughly speaking, given a f.p. $n$-parameter persistence module $\mathbb{M}$, an approximation parameter $\delta>0$, a $\delta$-grid of lines $L=L_{\delta}(K^{\delta})$ where $K$ compactly characterizes $\mathbb{M}$, and a matching function $\sigma$ (see Section 6 for a discussion about how to find such matching functions), Algorithm 1 works in three steps: